2-6(y+1)=4(2-3y)+6
2-6y+6=8-12y+6
-6y+12y=2+8+6
6y=16

Answer by mananth(16946)

2-6(y+1)=4(2-3y)+6
2-6y-6=8-12y+6
-6y+12y=-2+8+6
6y=12
divide by 6
y=2